OpenStudy (anonymous):

An athlete performing a long jump leaves the ground at a 29.6∘ angle and lands 7.68m away. Part A What was the takeoff speed? Express your answer using three significant figures and include the appropriate units. v =

OpenStudy (anonymous):

Part B If this speed were increased by just 3.0%, how much longer would the jump be? Express your answer using two significant figures and include the appropriate units. Δd =

OpenStudy (anonymous):

do we use that same formula here too

OpenStudy (anonymous):

Same thing as before:\[d=\frac{ v_0^2 }{ g }\sin(2\theta)\]You have the distance, you have the angle and you know g. the rest is easy

OpenStudy (anonymous):

7.69 = (sin (2 x 29.6) X 9.8 ) / 2

OpenStudy (anonymous):

wrong, review the formula

OpenStudy (anonymous):

sorry 7.69 = v0 (sin (2 x 29.6) / 9.8

OpenStudy (anonymous):

Also wrong, it is squared velocity

OpenStudy (anonymous):

The formula to get velocity is \[v_0=\sqrt{\frac{ d·g }{ \sin(2\theta)}}\]

OpenStudy (anonymous):

ohh whoops, 7.69 x 9.8 / sin (2 x 29.6) all squared root

OpenStudy (anonymous):

\[v_0=\sqrt{\frac{ 7.68·9.81 }{ \sin(2·29.6)}}\]

OpenStudy (anonymous):

i got an answer of 12.352

OpenStudy (anonymous):

For part number 2: Now you have a new speed that is a 3% faster than the initial speed\[v_1=1.03v_0\rightarrow d_1=\frac{ v_1^2 }{ g}\sin(2\theta)=1.03^2\left( \frac{ v_0^2 }{g} \sin(2\theta) \right)=1.03^2·d\]

OpenStudy (anonymous):

\[\Delta d=\frac{ d_1-d }{ d }=\frac{ (1.03^2-1)d }{ d }=1.03^2-1=0.06090\]And that means distance has increased in a 6.09%

OpenStudy (anonymous):

you got the 1.03 bc its 100% plus 3%, but how do you know to add it with 100%

OpenStudy (anonymous):

If the new velocity is 103% the initial velocity then you have to multiply by 103 and divide by 100, in other words, multiply by 1.03

OpenStudy (anonymous):

ohh what formula would i use for part b then? the same range formula?

OpenStudy (anonymous):

Please, read my explanation for part 2 above

OpenStudy (anonymous):

ohh okay i got the 3% part but i still dont understand the 2nd part, is this what you are trying to do... d= 1.03 sin(2 x 29.6) / 9.8 = 0.09

OpenStudy (anonymous):

Same formula and new speed. now speed is V1=1.03·V0 and that new speed will give a new distance d1

OpenStudy (anonymous):

so the speed we found in part a we use for this part b to get d=?

OpenStudy (anonymous):

but where do i use the 1.03 since that is a percentage

OpenStudy (anonymous):

Do you need to use it? dont you see that it is not necessary? Please, read what I have done carefully

OpenStudy (anonymous):

Ther are two posts where I describe in detail how to calculate what the problem is asking for in part 2. Please, come back to me when you read it carefully. You need to do some effort to try to understand what is written

OpenStudy (anonymous):

ohh so after finding d we multiply it by the 1.03 bc it increased in 3% ..right?

OpenStudy (anonymous):

Well let us leave it for today. You still do not want to read carefully what is written. Nobody has multiplied the distance by 1.03 Have a good night!

OpenStudy (anonymous):

i am looking at the formula and trying to understand it but its not sinking in

OpenStudy (anonymous):

@ganeshie8 do have another way of explaining part b, i am trying to figure this out and carlosGP helped me out but i still dont understand part b